A balloon is inflated with 0.35 moles of helium to a pressure of 1.63 atm. If the desired volume of the balloon is 1.15 L, what is the temperature in Kelvin? Round to two decimal places.
1
Expert's answer
2019-03-18T02:23:08-0400

P(He)×V(He) = n(He) ×R×T(He)

R = 0.082 L×atm×mol-1×K-1

T(He) = P(He) × V(He) / (n(He) ×R) = 1.63 atm × 1.15 L / (0.35 mol × 0.082 L×atm×mol-1×K-1) = 65.3136 K ≈ 65.31 K.